OpenStudy (anonymous):

What is the value of coefficient c in this quadratic equation? 5x^2 + 6x + 7 = 2(x + 2) 7 4 3 -3 3/5

OpenStudy (anonymous):

to have a quadratic equasion, you fist need to set one side = to 0

OpenStudy (anonymous):

subtract 2(x-2) from both sides

OpenStudy (anonymous):

5x^2 + 4x + 3 = 0 is what u get

OpenStudy (anonymous):

then think about what 'c' is in a quadratic equation

OpenStudy (anonymous):

a quatratic is ax^2 + bx + c = 0

OpenStudy (anonymous):

c in this equation is 3

OpenStudy (mrnood):

@grandzam btw it's subtract 2(x+2) from both sides

OpenStudy (anonymous):

eh that was a mistake. But I did the work right

OpenStudy (anonymous):

so it is 3? right? @MrNood

OpenStudy (mrnood):

@GracieFerrell Do you understand what @grandzam did above? You need to get the equation into the form ax^2+bx+c=0 if you can see how he did that , then you can see whether his answer is correct...

OpenStudy (anonymous):

yes I understand I was just checking